# Confocal Laser Endomicroscopy CLE-VS CLE video sequence

### v1

@@ 1 Introduction
One adjunct to the classical biopsy that has been shown in many recent studies to be used non-invasively and in-vivo in the anatomic region of the head and neck is Confocal Laser Endomicroscopy (CLE), an optical biopsy technique [1]. It delivers real-time diagnostic information about tissue structures at a highly magnified scale [1] and supports intraoperative margin assessment and surgical guidance [2]. The challenging aspect of CLE images is in interpreting the data, especially for untrained readers [2], where computer-assisted classification can be valuable in the intraoperative field. At present, CLE is not universally available and is routinely implemented only in a limited number of centers. Moreover, obtaining labeled data is challenging, as reliable correlates are hard to obtain given the margin-sampling constraints described above. For models with a vast amount of parameters, such as current deep learning models, this data shortage easily leads to overfitting, and hence to a severely reduced performance at inference. Few-shot learning methods, recently evaluated for CLE imaging [3], offer a potential solution, though classification with limited patient data remains challenging. Another recent strategy in this field is pretraining using self-supervised learning (SSL) methods, combined with adaptation to downstream tasks using limited datasets. Given the difficulty of obtaining histopathologic labels in CLE, exploiting unlabeled data becomes particularly valuable. This is especially relevant since ImageNet pretraining introduces a major domain shift, with features poorly capturing the distinct texture and contrast of CLE images.
@@ 3.1 Self Supervised Learning
SSL, i.e., training of models with a self-created supervisory signal, has recently emerged as a valuable tool in few-shot learning scenarios in medical imaging [4]. At its core is the idea to utilize large amounts of unlabeled data to pre-train feature extractors, which are subsequently used in downstream tasks using methods such as linear probing or adaptation. This setup helps to regularize the training process, which is key in few-shot scenarios. Caron et al. proposed self‑distillation of vision transformer models by training a teacher and a student on crops of the same image, encouraging the student to mimic the teacher’s predictions across views [5]. This yields robust, view‑invariant representations that generalize well to downstream few‑shot medical tasks, outperforming vanilla pre‑training and providing a simple yet powerful regularization that mitigates overfitting [5]. SSL was shown to significantly benefit from data deduplication [6], which can be attributed to eliminating conflicting signals in the loss function.
In this work, we train a vision transformer using the original DINO loss [5]. To avoid overfitting, allow for larger batch sizes, and expedite training, we chose the small configuration (ViT-small). For the training, we used AdamW as the optimizer and trained the model until convergence, as observed by the validation loss. We then retrospectively selected the model with the best validation loss in training.
@@ 3.2 CLE Video sequence filtering (CLE-ViFi)
A core contribution of this work is the inception of a video sequence filtering algorithm. A characteristic of CLE video sequences is their high inter-frame correlation, resulting from the relatively steady positioning of the CLE probe during surgical observation. Repeatedly showing nearly identical frames during training is inefficient and, more critically, can introduce conflicting supervisory signals and exacerbate dataset imbalance. As demonstrated by Oquab et al. [6], dataset variability is a key factor for successful SSL. In order to reduce the redundancy in the dataset, we propose a tailored video filtering scheme for CLE (CLE-ViFi).
@@
To support the SSIM kernel and reduce the sensor-noise within the CLE images, we compared different image scalings for their discriminatory power, and found the optimum for a scaling factor of 1/32. We found a sensible operating point for our application at a compromise between false negatives rate (0.1) and false positives (0.32) at a classification threshold $\tau$ of 0.411 (see Fig. 2 right).
Using this frame-based similarity classifier, we now process the entire HAN dataset. We select a first key frame and compare all subsequent frames until we reach a $SSIM<\tau$. This new frame is then added to the filtered dataset and selected as next key frame. We repeat this until we reach the end of the sequence. This filtering results in the HAN-Fi dataset, comprising 52,250 non-redundant CLE frames.
@@ 3.3 Downstream Task
For the downstream task, we employed SSL-pretrained models using both versions (HAN, HAN-Fi) of the dataset. We used linear probing, i.e., the training of a final classification layer on top of the SSL-trained feature extractor. We trained the network until convergence, using SGD with a momentum of 0.9 and an initial learning rate of 0.0001 with cosine annealing.
@@ 4 Results and Discussion
We furthermore can see a slight advantage in downstream task performance for the feature extractors trained on the filtered (HAN-Fi) dataset, leading to the conclusion that the video filtering was not detrimental to SSL training. However, by reducing the number of frames in SSL-based pretraining by approximately a factor of three, we significantly reduced the training time, as we found model convergence already after 2:27 RTX4090-GPU hours, compared to 7:23 GPU hours for the unfiltered dataset.
@@
ViT-small HNT [ours] linear probing $67.20\%\pm$33.83 $72.53\%\pm$13.19
HNT-ViFi [ours] linear probing $\textbf{67.48\%}\pm$34.28 $\textbf{73.52\%}\pm$12.52
Our investigations support the findings of the current studies in the field of AI research, which predominantly found that SSL-based pretraining can lead to improved results in specialized medical imaging tasks [7]. Specifically, our results show that SSL approaches enable more effective use of unlabeled data, which is a major advantage in a field with limited annotated image material. At the same time, our dataset filtering approach demonstrated a reduction in training time without any decrease in test accuracy across both downstream datasets. Our results show considerable fluctuations, as can be seen in the high standard deviations in Tab. 1. This stems from the substantial heterogeneity of the datasets rather than from training instability: the test set, in particular, includes patients with highly diverse and complex anatomical structures, leading to the observed high standard deviations.
@@
The authors would like to acknowledge support by the German Research Foundation (DFG), project number 545049923.

### v2

@@ 1 Introduction
One adjunct to the classical biopsy that has been shown in many recent studies to be used non-invasively and in-vivo in the anatomic region of the head and neck is Confocal Laser Endomicroscopy (CLE), an optical biopsy technique [sievert-2021-review-cle-hnsscc]. It delivers real-time diagnostic information about tissue structures at a highly magnified scale [sievert-2021-review-cle-hnsscc] and supports intraoperative margin assessment and surgical guidance [VILLARD2022105826]. The challenging aspect of CLE images is in interpreting the data, especially for untrained readers [VILLARD2022105826], where computer-assisted classification can be valuable in the intraoperative field. At present, CLE is not universally available and is routinely implemented only in a limited number of centers. Moreover, obtaining labeled data is challenging, as reliable correlates are hard to obtain given the margin-sampling constraints described above. For models with a vast amount of parameters, such as current deep learning models, this data shortage easily leads to overfitting, and hence to a severely reduced performance at inference. Few-shot learning methods, recently evaluated for CLE imaging [aubreville2023-fewshotlearningcle], offer a potential solution, though classification with limited patient data remains challenging. Another recent strategy in this field is pretraining using self-supervised learning (SSL) methods, combined with adaptation to downstream tasks using limited datasets. Given the difficulty of obtaining histopathologic labels in CLE, exploiting unlabeled data becomes particularly valuable. This is especially relevant since ImageNet pretraining introduces a major domain shift, with features poorly capturing the distinct texture and contrast of CLE images.
@@ 3.1 Self Supervised Learning
SSL, i.e., training of models with a self-created supervisory signal, has recently emerged as a valuable tool in few-shot learning scenarios in medical imaging [ouyang2022self]. At its core is the idea to utilize large amounts of unlabeled data to pre-train feature extractors, which are subsequently used in downstream tasks using methods such as linear probing or adaptation. This setup helps to regularize the training process, which is key in few-shot scenarios. Caron et al. proposed self‑distillation of vision transformer models by training a teacher and a student on crops of the same image, encouraging the student to mimic the teacher’s predictions across views [caron2021emerging]. This yields robust, view‑invariant representations that generalize well to downstream few‑shot medical tasks, outperforming vanilla pre‑training and providing a simple yet powerful regularization that mitigates overfitting [caron2021emerging]. SSL was shown to significantly benefit from data deduplication [oquab2023dinov2], which can be attributed to eliminating conflicting signals in the loss function.
In this work, we train a vision transformer using the original DINO loss [caron2021emerging]. To avoid overfitting, allow for larger batch sizes, and expedite training, we chose the small configuration (ViT-small). For the training, we used AdamW as the optimizer and trained the model until convergence, as observed by the validation loss. We then retrospectively selected the model with the best validation loss in training.
@@ 3.2 CLE Video sequence filtering (CLE-ViFi)
A core contribution of this work is the inception of a video sequence filtering algorithm. A characteristic of CLE video sequences is their high inter-frame correlation, resulting from the relatively steady positioning of the CLE probe during surgical observation. Repeatedly showing nearly identical frames during training is inefficient and, more critically, can introduce conflicting supervisory signals and exacerbate dataset imbalance. As demonstrated by Oquab et al. [oquab2023dinov2], dataset variability is a key factor for successful SSL. In order to reduce the redundancy in the dataset, we propose a tailored video filtering scheme for CLE (CLE-ViFi).
@@
To support the SSIM kernel and reduce sensor-noise within the CLE images, we compared different image scalings for their discriminatory power and found the optimum scaling factor to be 1/32. We found a sensible operating point for our application at a compromise between false negatives rate (0.1) and false positives (0.32) at a classification threshold $\tau$ of 0.411 (see Fig. 2 right).
Using this frame-based similarity classifier, we now process the entire HAN dataset. We select a first key frame and compare all subsequent frames until we reach a $SSIM<\tau$. This new frame is then added to the filtered dataset and selected as the next key frame. We repeat this until we reach the end of the sequence. This filtering results in the HAN-ViFi dataset, comprising 52,250 non-redundant CLE frames.
@@ 3.3 Downstream Task
For the downstream task, we employed SSL-pretrained models using both versions (HAN, HAN-ViFi) of the dataset. We used linear probing, i.e., the training of a final classification layer on top of the SSL-trained feature extractor. We trained the network until convergence, using SGD with a momentum of 0.9 and an initial learning rate of 0.0001 with cosine annealing.
@@ 4 Results and Discussion
We furthermore can see a slight advantage in downstream task performance for the feature extractors trained on the filtered (HAN-ViFi) dataset, leading to the conclusion that the video filtering was not detrimental to SSL training. However, by reducing the number of frames in SSL-based pretraining by approximately a factor of three, we significantly reduced the training time, as we found model convergence already after 2:27 RTX4090-GPU hours training iterations, compared to 7:23 GPU hours for the unfiltered dataset.
@@
ViT-small HAN [ours] linear probing $67.20\%\pm$33.83 $72.53\%\pm$13.19
HAN-ViFi [ours] linear probing $\textbf{67.48\%}\pm$34.28 $\textbf{73.52\%}\pm$12.52
Our investigations support the findings of the current studies in the field of AI research, which predominantly found that SSL-based pretraining can lead to improved results in specialized medical imaging tasks [huang2023self]. Specifically, our results show that SSL approaches enable more effective use of unlabeled data, which is a major advantage in a field with limited annotated image material. At the same time, our dataset filtering approach demonstrated a reduction in training time without any decrease in test accuracy across both downstream datasets. Our results show considerable fluctuations, as can be seen in the high standard deviations in Tab. 1. This stems from the substantial heterogeneity of the datasets rather than from training instability: the test set, in particular, includes patients with highly diverse and complex anatomical structures, leading to the observed high standard deviations.
@@
Withheld for blind peer review. The authors would like to acknowledge support by the German Research Foundation (DFG), project number 545049923.